OpenStudy (anonymous):

What is the solution to the following system using Cramer's Rule? 3x – 6y = –1 5x + 3y = 7 (1, 3) (–1, 3) (–1,2/3 ) (1, 2/3)

OpenStudy (amistre64):

well, the solution to the system will be the same regardless of the method; to just get the correct option, plug them into the equations to see what works out

OpenStudy (anonymous):

i tried, but i didnt get anything that worked.

OpenStudy (amistre64):

3x – 6y = –1 1 3 3 - 18 .... not it 5x + 3y = 7 -1 3 -5 + 9 .... not it

OpenStudy (anonymous):

(1,2/3)Via Mathematica:\[\text{LinearSolve}\left[\left( \begin{array}{cc} 3 & -6 \\ 5 & 3 \end{array} \right),\left( \begin{array}{c} -1 \\ 7 \end{array} \right)\right]=\left\{1,\frac{2}{3}\right\} \]

OpenStudy (amistre64):

3x – 6y = –1 -1 2/3 -3 -4 ... not it itll have to be the last one 5x + 3y = 7

OpenStudy (anonymous):

shana, to use cramers rule.. set up your matrix

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

3x – 6y = –1 1 2/3 3 - 4 = -1 5x + 3y = 7 1 2/3 5 2 = 7

OpenStudy (amistre64):

yep